### v1

@@ Abstract
Additionally, the fidelity of the widely used XY8 sequence Souza et al. (2012) with phases $\left(0,1,0,1,1,0,1,0\right)\pi/2$ is
@@
Usually the transition probability error is quite small, i.e., $\epsilon\ll 1$, e.g., $\epsilon=1-\cos^{2}\left(\Delta t/2\right)\approx 0.006$ for a target $\omega_{L}t=\pi$ pulse and $\Delta/\omega_{L}=0.05$, assuming $\delta=0$, so the error in the fidelity ($1-F$) of the XY8 sequence ($\sim\epsilon^{3}$) will be much smaller than the one of the sequence with constant zero phases ($\sim\epsilon$). A numerical simulation of the fidelity of the continuous (labeled CX) drive and the XY8 sequence with zero pulse separation (labeled CXY8) is shown in Fig. S.1. The simulation shows that CXY8 is significantly more robust than the standard continous decoupling. Similarly, one can show that we can obtain a robust performance and even better fidelity with other sequences of phased pulses, e.g., by using the KDD or UR sequences Souza et al. (2012); Genov et al. (2014, 2017); Bruns et al. (2018).
Finally, we demonstrate the robustness of CPDD for sensing the amplitude $g\approx 2\pi~{}11.7$ kHz of a weak oscillating field. The results from a numerical simulation and experiments are shown in Fig. 2 in the main text and the details for the simulation results are demonstrated in Fig. S.2. Figure S.2 shows that when the Rabi frequency of the driving field is equal to the frequency of the sensed field, i.e., $\Omega=\omega_{L}=2\pi~{}8$ MHz, we observe oscillations of the population of state $|0\rangle$ for both CX and CXY8 with $P=\cos(\Theta(t)/2)$, where $\Theta=gt$. The oscillations are quickly lost with CX even with small $\Delta\approx 2\pi~{}20$ kHz, i.e., of the order of the coupling strength $g$. On the contrary, with CXY8 the robustness is improved by more than twenty times, as we observe oscillations even for $\Delta=2\pi~{}400$ kHz Sup .
@@
In this section, we analyze how CPDD can be used for the detection of a weak oscillating signal. We consider a measurement, which consists of the microwave pulse sequence $\pi/2(x)$ pulse - dynamical decoupling (DD) - $\pi/2(\phi)$ pulse, where DD can be pulsed, continuous, or CPDD and the phase $\phi$ of the last $\pi/2$ pulse can be varied. If we apply the dynamical decoupling sequence on resonance, we observe Rabi oscillations, characterized with a pulse area $\Theta$, which can be observed directly in the bare basis, as evident from Eq. (S.1.1) and Eq. (S.10). On resonance, the pulse area $\Theta=2\alpha gt\cos{\left(\xi\right)}$, where the attenuation factor $\alpha=2/\pi$ for pulsed dynamical decoupling and $\alpha=1/2$ for CPDD. In the following, we consider the application of CPDD for slope/linear and variance detection Degen et al. (2017) and compare its performance to standard techniques.
@@
Slope/linear detection is usually applicable when the initial phase $\xi$ of the sensed field, e.g., $B\cos{\left(\omega_{s}t+\xi\right)}$, is fixed at the beginning of each measurement Degen et al. (2017). The measurement protocol for determining the amplitude $B$ is as described above with the phase of the last $\pi/2$ pulse alternating between $90^{\circ}$ and $-90^{\circ}$, i.e., we apply $\pi/2(\pm y)$ pulses. The population of state $|0\rangle$ of the NV electron spin with the last $\pi/2(\pm y)$ pulses is then $p_{0}=\frac{1}{2}\mp\frac{1}{2}\sin\left(\Theta\right)$. The normalized signal contrast from the fluorescence difference from the two measurements is then Staudenmaier et al. (2022, 2023)
@@
$=c_{\text{max}}e^{-\Gamma(t)}\sin\left(\alpha\gamma_{\text{NV}}Bt\cos{\left(\xi\right)}\right)$
$\approx c_{\text{max}}e^{-\Gamma(t)}\alpha\gamma_{\text{NV}}Bt\cos{\left(\xi\right)}$ (S.19)
where $c_{\text{max}}$ is the maximum contrast, $\Gamma(t)$ accounts for decoherence during the duration $t$ of the decoupling sequence, and $\Theta=2\alpha gt\cos{\left(\xi\right)}=\alpha\gamma_{\text{NV}}Bt\cos{\left(\xi\right)}$, as $g=\gamma_{\text{NV}}B/2$ in our experiment. The last approximation is valid only for small $\Theta\ll 1$. One can estimate the amplitude of the oscillating magnetic field by fitting the observed Rabi oscillations, similar to the experiment in Fig. 2 in the main text with $\Theta=\alpha\gamma_{\text{NV}}Bt\cos{\left(\xi\right)}$.
Another approach is to detect a change in the signal from a reference point of the measurement, which corresponds to a well-known value of $B$. Slope or linear detection is typically used to detect a perturbation in a signal with a constant initial phase Degen et al. (2017). The optimal reference point is when $p_{0}=1/2$ Degen et al. (2017), i.e., $\Theta=\pi k,k\in\mathbb{Z}$. Then, the expected change in the contrast $dc$ due to a small perturbation $dB$ is given by
@@
$=\mp c_{\text{max}}e^{-\Gamma(t)}\alpha\gamma_{\text{NV}}t\cos{\left(\xi\right)}\,dB,$ (S.20)
where the chosen measurement time duration is typically close to $t=T_{2}/2$ for the best sensitivity, given a typical decay function $\Gamma(t)$ with coherence time $T_{2}$ Degen et al. (2017). It is evident that the detection mechanism of CPDD is similar to the one of the standard pulsed techniques with the main difference being the attenuation factor $\alpha$.
@@
When the initial phase and/or amplitude of the detected oscillating signal fluctuate from measurement to measurement it is advantageous to detect the signal variance. We then consider a measurement, which consists of the microwave pulse sequence $\pi/2(x)$ pulse - dynamical decoupling (DD) - $\pi/2(\pm x)$ pulse, where DD can be pulsed, continuous, or CPDD. If we apply the respective sequence on resonance, we observe Rabi oscillations with a pulse area $\Theta$, which can be observed directly in the bare basis, as evident from Eq. (S.1.1) and Eq. (S.10). The population of state $|0\rangle$ of the NV electron spin with the last $\pi/2(\pm x)$ pulses is then $p_{0}=\frac{1}{2}\mp\frac{1}{2}\cos\left(\Theta\right)$. The normalized signal contrast from the fluorescence difference from the two measurements is Staudenmaier et al. (2022)
@@
$\overline{c}$ $=c_{\text{max}}e^{-\Gamma(t)}\left\langle\cos\left(\alpha\gamma_{\text{NV}}Bt\cos{\left(\xi\right)}\right)\right\rangle$ (S.23)
$=c_{\text{max}}e^{-\Gamma(t)}J_{0}\left(\alpha\gamma_{\text{NV}}Bt\right)$
$\approx c_{\text{max}}e^{-\Gamma(t)}(1-\alpha^{2}\gamma_{\text{NV}}^{2}B^{2}t^{2}/4),$
$=c_{\text{max}}e^{-\Gamma(t)}(1-\alpha^{2}\gamma_{\text{NV}}^{2}B_{\text{rms}}^{2}t^{2}/2),$
@@
Next, we consider the case when both the amplitude and the phase of the signal vary from experiment to experiment. For example, this is the case when the proton signal we observe is due to statistical polarization in the detection volume of the NV center. We assume that the magnetic field, generated by the sample due to statistical polarization is given by Gefen et al. (2019); Staudenmaier et al. (2022)
@@
$=c_{\text{max}}e^{-\Gamma(t)}\exp\left(-\alpha^{2}\gamma_{\text{NV}}^{2}B_{\text{rms}}^{2}t^{2}/2\right)$
$\approx c_{\text{max}}e^{-\Gamma(t)}\left(1-\alpha^{2}\gamma_{\text{NV}}^{2}B_{\text{rms}}^{2}t^{2}/2\right),$ (S.25)
where we used that $\langle\Theta^{2}\rangle\equiv\Theta_{\text{rms}}^{2}=\alpha^{2}\gamma_{\text{NV}}^{2}\langle B^{2}\rangle t^{2}$ with $\langle B^{2}\rangle=B_{\text{rms}}^{2}$. The last approximation is valid only for small $\langle\Theta^{2}\rangle$. We detect $B_{\text{rms}}$ by fitting the experimental values to the decay function in Eq. (S.3.2) when we vary the duration $t$ of the decoupling sequence. Similar to linear detection, the performance of CPDD for variance detection is similar to the one of the standard pulsed techniques with the main difference being the attenuation factor $\alpha$.
Figure S.3: Simulated spectra showing the population of state $|0\rangle$ as a function of the detection (angular) frequency of CPDD $\pi/T$ in units of $\omega_{L}$ with phase changes at time intervals $T$, which varied between 100 and 400 ns with the Rabi frequency $\Omega=\pi/T$. The sensed field has an angular frequency $\omega_{L}=2\pi~{}2$ MHz, an amplitude of $g=2\pi~{}20$ kHz, and initial phase $\xi=0$. The ratio $g/\Omega=0.01$ on resonance and the total number of $480$ pulses correspond to the respective parameters in simulations of spurious harmonics with pulsed DD in Loretz et al. (2015). The sensor qubit is initially prepared in state $|0\rangle$, e.g., by optical pumping. CPDD with the CXY4 sequence shows the expected spurious harmonics at $32\omega_{L}/$ 3 and $2\omega_{L}$, similar to pulsed DD with XY4 Loretz et al. (2015). CPDD with the CXY4R, which uses phase randomization, allows one to suppress these harmonics almost completely, similar to pulsed DD Wang et al. (2019). The simulation with randomized phases is performed by adding a random phase to each of the $120$ sequences in a single experiment and running the simulation $640$ times. The incomplete suppression around $32\omega_{L}/$ 3 is most likely due to the limited number of runs and the higher ratio $g/\Omega$.
@@
Similar to pulsed dynamical decoupling Loretz et al. (2015), the effect of the signal during sensing can produce spurious harmonics, which depend on the applied sequence of phase changes (see Fig. S.3). Since the signal is sensed by its effect during the sequence, the spurious harmonics can be comparable to the main one. This is evident in the simulation in Fig. S.3, where the black curve shows the expected spurious harmonics with CPDD with the CXY4 sequence at $32\omega_{L}/$ 3 and $2\omega_{L}$, in a similar way to pulsed dynamical decoupling with XY4 Loretz et al. (2015). CPDD with the CXY4R, which uses phase randomization, allows one to suppress these harmonics almost completely, in a similar way to pulsed DD Wang et al. (2019). The simulation parameters correspond to the respective spurious harmonics with pulsed DD in Ref. Loretz et al. (2015).
@@
Two diamond samples where used for the experiments. For the measurements shown Figure 3 of the main text the diamond sample is an electronic grade, type IIa synthetic diamond featuring an isotopically enriched 12C layer of $\approx 250~{}\mathrm{\mu m}$ thickness. The layer with an isotopic purity of $9$ was grown by Element Six using chemical vapor deposition. To create shallow NV centers suitable for the detection of hydrogen spins on the diamond surface, ion implantation was employed using nitrogen ions with 98 % 15N isotopic abundance. An ion energy of $2.5~{}\mathrm{keV}$ was chosen to yield NV centers at typical depths of $8\pm 2~{}\mathrm{nm}$ Findler et al. (2020). A nitrogen ion dose of $5\times 10^{8}~{}\frac{{}^{15}N^{+}}{cm^{2}}$ results in single, isolated NV centers. After ion implantation the sample was annealed in ultra high vacuum with pressures below $1\times 10^{-7}~{}\mathrm{mbar}$ at $1000~{}\mathrm{{}^{\circ}C}$ for 3 hours to mobilize the created vacancies and allow for NV center creation with a yield of 0.7 %. Before each processing step, the sample was boiled several hours in a 1:1:1 mixture of sulfuric, perchloric and nitric acid to remove any graphitic or organic residues from the surface. The NV is about $d=(10.1\pm 0.5)$ nm deep in the diamond, which we estimated by measuring the variance of the fluctuating NMR magnetic field at its location with XY8 dynamical decoupling Müller et al. (2014).
@@
Figure S.4: (a) Pulse scheme for the CXY8-$N$ and XY8-$N$ measurements for frequency sensing. We use a laser to prepare the electron spin of an NV center in state $|0\rangle$ by optical pumping and read out its final state. We prepare a coherent superposition state by a microwave $\pi/2$ pulse, apply the control scheme (CXY8 or XY8), and then another $\pi/2$ pulse with a phase that alternates, to map coherences back onto populations. The final population of state $|0\rangle$ is estimated from the difference of the signals from the two alternating measurements, reducing errors due to charge-state and count-rate fluctuations Häberle et al. (2017).We vary $T$ ($\tau$) for CXY8-$N$ (XY8-$N$) with fixed repetition number $N$ to detect the frequency of the AC signal. (b) Measured signal, corresponding to the population of state $|0\rangle$ of the electron spin of a single NV center, vs. the detected signal frequency $f=\nicefrac{{1}}{{2T}}$ ($f=\nicefrac{{1}}{{2\tau}}$) for CXY8 (XY8). We observe a dip at the frequency 170017001700\,1700kHz of the artificial signal for both CXY8-5 and XY8-5 sequences, as expected from theory. The slightly lower population with XY8 than CXY8 on resonance is expected and due to the difference in the rectified signal amplitudes of the two sequences (see text). The vertical dashed lines for all subfigures show the expected full width at half maximum (FWHM) of the dip, i.e., the inverse of the sequence duration, which is FWHM $\approx 85\,$kHz in this case. (c) We observe the expected a dip in the population at 800080008000\,8000kHz of a high frequency artificial signal, where only CXY8 is applicable. FWHM $\approx 83.3\,$kHz. (d) The measurement is the same as in (b) but with a lower artificial signal amplitude, which requires a higher sequence order and total sequence duration to observe a similar dip; FWHM $\approx 16.3\,$kHz. The additional sideband dip with XY8 does not correspond to the frequency of a real signal and is likely due to pulse errors of the non-instantaneous rectangular pulses. (e) The measurement is same as in (d) but with a lower artificial signal amplitude; FWHM $\approx 16.4\,$kHz due to the slightly shorter sequence duration than in (d). All experimental results show excellent correspondence to theory.
@@
All MW pulses for the experimental sequences were generated in direct synthesis by a Keysight (M8195A) and Tektronix (AWG70001A) arbitrary waveform generators, using a Python based in house written software (Qudi) Binder et al. (2017). The measurement sequences are programmed to keep the phase relation between all pulses in the rotating frame and the high sampling rate of the AWGs (50 and 60 GS/s) allows to have precise direct control of the signal phases. The instantaneous rise time of the devices allows for a bandwidth as the inverse of twice the sampling rate.
@@
In a first series of experiments, we use a single NV center to detect the frequency of an artificially generated AC signal at frequency 170017001700\,1700 kHz (800080008000\,8000 kHz), which we label low (high) frequency for two different signal amplitudes. Figure S.4(a) shows the pulse scheme for the CXY8-$N$ and XY8-$N$ measurements. We use a laser to prepare the electron spin of the NV center in state $|0\rangle$ by optical pumping and read out its final state. We prepare a coherent superposition state by a microwave $\pi/2$ pulse, apply the control scheme (CXY8 or XY8), and then use another $\pi/2$ pulse with a phase that alternates between 00 and $180^{\circ}$, to map coherences back onto populations. We estimate the final population of state $|0\rangle$ from the difference between the signals, reducing errors due to charge-state and count-rate fluctuations Häberle et al. (2017). All control fields originate from the same arbitrary waveform generator and amplifier. We generate the artificial external signal to be sensed from a second microwave source, so the phase of the signal at the beginning of the interaction is not fixed with respect to the control fields. We vary the interval of phase changes $T$ (the interpulse duration $\tau$) for CXY8 (XY8) to detect the frequency of the AC signal, using the relation $f=\nicefrac{{1}}{{2T}}$ ($f=\nicefrac{{1}}{{2\tau}}$).
Figure S.4(b) demonstrates frequency sensing of a relatively low frequency (1700 kHz) AC signal, which is typical for pulsed DD applications for NV centers. As expected from theory, the measured signal shows a dip in the population of state $|0\rangle$ at the frequency 170017001700\,1700kHz of the artificial signal, for both CXY8-5 and XY8-5 sequences. The slightly lower population with XY8 than CXY8 on resonance is expected from theory and is due to the corresponding difference in the rectified signal amplitudes of the two sequences, which are $g/2$ ($2g/\pi$) for CXY8 (XY8) with $g$ – the amplitude of the artificial signal in angular frequency units (see the main text). The expected full width at half maximum (FWHM) of the dip also matches theory and is approximately equal to the inverse of the sequence duration $8N\times\nicefrac{{1}}{{2f}}$ ms on resonance. Thus, FWHM $\approx\nicefrac{{2f}}{{8N}}=\nicefrac{{2\times 1700}}{{8\times 5}}=85\,$kHz for both CXY8 and XY8.
Figure S.4(c) demonstrates the application of CXY8-24 for detecting a high frequency AC field, where standard pulsed DD, e.g., with the XY8 sequence, cannot be applied. We observe the expected dip in the population at the frequency 800080008000\,8000kHz of the artificial signal with a FWHM $\approx\nicefrac{{2f}}{{8N}}=\nicefrac{{2\times 8000}}{{8\times 24}}=83.3\,$kHz.
@@
We also perform quantum heterodyne (Qdyne) measurements with the single NV center and an 8 MHz test signal. For details on the measurement technique see the main text and references Schmitt et al. (2017); Staudenmaier et al. (2021, 2023).
@@
Figure 4 in the main text shows the measured spectrum of the Qdyne measurement. In Figure S.6 we show further analysis of the performance of the Qdyne measurement. The analysis is done equivalently to Ref. Staudenmaier et al. (2021). The total measurement time is $T_{\mathrm{tot}}=120$ s. For the analysis we slice the data into smaller parts. The shortest time is 0.57 s (216 samples). For every measurement time the FFT of the corresponding time trace is calculated. A Lorentzian fit of the form

### v2

@@ Abstract
Additionally, the fidelity of the widely used XY8 sequence [36] with phases $\left(0,1,0,1,1,0,1,0\right)\pi/2$ is
@@
Usually the transition probability error is quite small, i.e., $\epsilon\ll 1$, e.g., $\epsilon=1-\cos^{2}\left(\Delta t/2\right)\approx 0.006$ for a target $\omega_{L}t=\pi$ pulse and $\Delta/\omega_{L}=0.05$, assuming $\delta=0$, so the error in the fidelity ($1-F$) of the XY8 sequence ($\sim\epsilon^{3}$) will be much smaller than the one of the sequence with constant zero phases ($\sim\epsilon$). A numerical simulation of the fidelity of the continuous (labeled CX) drive and the XY8 sequence with zero pulse separation (labeled CXY8) is shown in Fig. S.1. The simulation shows that CXY8 is significantly more robust than the standard continous decoupling. Similarly, one can show that we can obtain a robust performance and even better fidelity with other sequences of phased pulses, e.g., by using the KDD or UR sequences [36, 35, 38, 45].
Finally, we demonstrate the robustness of CPDD for sensing the amplitude $g\approx 2\pi~{}11.7$ kHz of a weak oscillating field. The results from a numerical simulation and experiments are shown in Fig. 2 in the main text and the details for the simulation results are demonstrated in Fig. S.2. Figure S.2 shows that when the Rabi frequency of the driving field is equal to the frequency of the sensed field, i.e., $\Omega=\omega_{L}=2\pi~{}8$ MHz, we observe oscillations of the population of state $|0\rangle$ for both CX and CXY8 with $P=\cos(\Theta(t)/2)$, where $\Theta=gt$. The oscillations are quickly lost with CX even with small $\Delta\approx 2\pi~{}20$ kHz, i.e., of the order of the coupling strength $g$. On the contrary, with CXY8 the robustness is improved by more than twenty times, as we observe oscillations even for $\Delta=2\pi~{}400$ kHz.
@@
In this section, we analyze how CPDD can be used for the detection of a weak oscillating signal. We consider a measurement, which consists of the microwave pulse sequence $\pi/2(x)$ pulse - dynamical decoupling (DD) - $\pi/2(\phi)$ pulse, where DD can be pulsed, continuous, or CPDD and the phase $\phi$ of the last $\pi/2$ pulse can be varied. If we apply the dynamical decoupling sequence on resonance, we observe Rabi oscillations, characterized with a pulse area $\Theta$, which can be observed directly in the bare basis, as evident from Eq. (S.1.1) and Eq. (S.10). On resonance, the pulse area $\Theta=2\kappa gt\cos{\left(\xi\right)}$, where the attenuation factor $\kappa=2/\pi$ for pulsed dynamical decoupling and $\kappa=1/2$ for CPDD [6]. In the following, we consider the application of CPDD for slope/linear and variance detection [6] and compare its performance to standard techniques.
@@
Slope/linear detection is usually applicable when the initial phase $\xi$ of the sensed field, e.g., $B\cos{\left(\omega_{s}t+\xi\right)}$, is fixed at the beginning of each measurement [6]. The measurement protocol for determining the amplitude $B$ is as described above with the phase of the last $\pi/2$ pulse alternating between $90^{\circ}$ and $-90^{\circ}$, i.e., we apply $\pi/2(\pm y)$ pulses. The population of state $|0\rangle$ of the NV electron spin with the last $\pi/2(\pm y)$ pulses is then $p_{0}=\frac{1}{2}\mp\frac{1}{2}\sin\left(\Theta\right)$. The normalized signal contrast from the fluorescence difference from the two measurements is then [60, 48]
@@
$=c_{\text{max}}e^{-\Gamma(t)}\sin\left(\kappa\gamma_{\text{NV}}Bt\cos{\left(\xi\right)}\right)$
$\approx c_{\text{max}}e^{-\Gamma(t)}\kappa\gamma_{\text{NV}}Bt\cos{\left(\xi\right)},$ (S.19)
where $c_{\text{max}}$ is the maximum contrast, $\Gamma(t)$ accounts for decoherence during the duration $t$ of the decoupling sequence, and $\Theta=2\kappa gt\cos{\left(\xi\right)}=\kappa\gamma_{\text{NV}}Bt\cos{\left(\xi\right)}$, as $g=\gamma_{\text{NV}}B/2$ in our experiment. The last approximation is valid only for small $\Theta\ll 1$. One can estimate the amplitude of the oscillating magnetic field by fitting the observed Rabi oscillations, similar to the experiment in Fig. 2 in the main text with $\Theta=\kappa\gamma_{\text{NV}}Bt\cos{\left(\xi\right)}$.
Another approach is to detect a change in the signal from a reference point of the measurement, which corresponds to a well-known value of $B$. Slope or linear detection is typically used to detect a perturbation in a signal with a constant initial phase [6]. The optimal reference point is when $p_{0}=1/2$ [6], i.e., $\Theta=\pi k,k\in\mathbb{Z}$. Then, the expected change in the contrast $dc$ due to a small perturbation $dB$ is given by
@@
$=\mp c_{\text{max}}e^{-\Gamma(t)}\kappa\gamma_{\text{NV}}t\cos{\left(\xi\right)}\,dB,$ (S.20)
where the chosen measurement time duration is typically close to $t=T_{2}/2$ for the best sensitivity, given a typical decay function $\Gamma(t)$ with coherence time $T_{2}$ [6]. It is evident that the detection mechanism of CPDD is similar to the one of the standard pulsed techniques with the main difference being the attenuation factor $\kappa$.
@@
When the initial phase and/or amplitude of the detected oscillating signal fluctuate from measurement to measurement it is advantageous to detect the signal variance. We then consider a measurement, which consists of the microwave pulse sequence $\pi/2(x)$ pulse - dynamical decoupling (DD) - $\pi/2(\pm x)$ pulse, where DD can be pulsed, continuous, or CPDD. If we apply the respective sequence on resonance, we observe Rabi oscillations with a pulse area $\Theta$, which can be observed directly in the bare basis, as evident from Eq. (S.1.1) and Eq. (S.10). The population of state $|0\rangle$ of the NV electron spin with the last $\pi/2(\pm x)$ pulses is then $p_{0}=\frac{1}{2}\mp\frac{1}{2}\cos\left(\Theta\right)$. The normalized signal contrast from the fluorescence difference from the two measurements is [60]
@@
$\overline{c}$ $=c_{\text{max}}e^{-\Gamma(t)}\left\langle\cos\left(\kappa\gamma_{\text{NV}}Bt\cos{\left(\xi\right)}\right)\right\rangle$ (S.23)
$=c_{\text{max}}e^{-\Gamma(t)}J_{0}\left(\kappa\gamma_{\text{NV}}Bt\right)$
$\approx c_{\text{max}}e^{-\Gamma(t)}(1-\kappa^{2}\gamma_{\text{NV}}^{2}B^{2}t^{2}/4),$
$=c_{\text{max}}e^{-\Gamma(t)}(1-\kappa^{2}\gamma_{\text{NV}}^{2}B_{\text{rms}}^{2}t^{2}/2),$
@@
Next, we consider the case when both the amplitude and the phase of the signal vary from experiment to experiment. For example, this is the case when the proton signal we observe is due to statistical polarization in the detection volume of the NV center. We assume that the magnetic field, generated by the sample due to statistical polarization is given by [50, 60]
@@
$=c_{\text{max}}e^{-\Gamma(t)}\exp\left(-\kappa^{2}\gamma_{\text{NV}}^{2}B_{\text{rms}}^{2}t^{2}/2\right)$
$\approx c_{\text{max}}e^{-\Gamma(t)}\left(1-\kappa^{2}\gamma_{\text{NV}}^{2}B_{\text{rms}}^{2}t^{2}/2\right),$ (S.25)
where we used that $\langle\Theta^{2}\rangle\equiv\Theta_{\text{rms}}^{2}=\kappa^{2}\gamma_{\text{NV}}^{2}\langle B^{2}\rangle t^{2}$ with $\langle B^{2}\rangle=B_{\text{rms}}^{2}$. The last approximation is valid only for small $\langle\Theta^{2}\rangle$. We detect $B_{\text{rms}}$ by fitting the experimental values to the decay function in Eq. (S.3.2) when we vary the duration $t$ of the decoupling sequence. Similar to linear detection, the performance of CPDD for variance detection is similar to the one of the standard pulsed techniques with the main difference being the attenuation factor $\kappa$.
Figure S.3: Simulated spectra showing the population of state $|0\rangle$ as a function of the detection (angular) frequency of CPDD $\pi/T$ in units of $\omega_{L}$ with phase changes at time intervals $T$, which varied between 100 and 400 ns with the Rabi frequency $\Omega=\pi/T$. The sensed field has an angular frequency $\omega_{L}=2\pi~{}2$ MHz, an amplitude of $g=2\pi~{}20$ kHz, and initial phase $\xi=0$. The ratio $g/\Omega=0.01$ on resonance and the total number of $480$ pulses correspond to the respective parameters in simulations of spurious harmonics with pulsed DD in [3]. The sensor qubit is initially prepared in state $|0\rangle$, e.g., by optical pumping. CPDD with the CXY4 sequence shows the expected spurious harmonics at $32\omega_{L}/$ 3 and $2\omega_{L}$, similar to pulsed DD with XY4 [3]. CPDD with the CXY4R, which uses phase randomization, allows one to suppress these harmonics almost completely, similar to pulsed DD [53]. The simulation with randomized phases is performed by adding a random phase to each of the $120$ sequences in a single experiment and running the simulation $640$ times. The incomplete suppression around $32\omega_{L}/$ 3 is most likely due to the limited number of runs and the higher ratio $g/\Omega$.
@@
Similar to pulsed dynamical decoupling [3], the effect of the signal during sensing can produce spurious harmonics, which depend on the applied sequence of phase changes (see Fig. S.3). Since the signal is sensed by its effect during the sequence, the spurious harmonics can be comparable to the main one. This is evident in the simulation in Fig. S.3, where the black curve shows the expected spurious harmonics with CPDD with the CXY4 sequence at $32\omega_{L}/$ 3 and $2\omega_{L}$, in a similar way to pulsed dynamical decoupling with XY4 [3]. CPDD with the CXY4R, which uses phase randomization, allows one to suppress these harmonics almost completely, in a similar way to pulsed DD [53]. The simulation parameters correspond to the respective spurious harmonics with pulsed DD in Ref. [3].
@@
Two diamond samples where used for the experiments. For the measurements shown Figure 3 of the main text the diamond sample is an electronic grade, type IIa synthetic diamond featuring an isotopically enriched 12C layer of $\approx 250~{}\mathrm{\mu m}$ thickness. The layer with an isotopic purity of $9$ was grown by Element Six using chemical vapor deposition. To create shallow NV centers suitable for the detection of hydrogen spins on the diamond surface, ion implantation was employed using nitrogen ions with 98 % 15N isotopic abundance. An ion energy of $2.5~{}\mathrm{keV}$ was chosen to yield NV centers at typical depths of $8\pm 2~{}\mathrm{nm}$ [46]. A nitrogen ion dose of $5\times 10^{8}~{}\frac{{}^{15}N^{+}}{cm^{2}}$ results in single, isolated NV centers. After ion implantation the sample was annealed in ultra high vacuum with pressures below $1\times 10^{-7}~{}\mathrm{mbar}$ at $1000~{}\mathrm{{}^{\circ}C}$ for 3 hours to mobilize the created vacancies and allow for NV center creation with a yield of 0.7 %. Before each processing step, the sample was boiled several hours in a 1:1:1 mixture of sulfuric, perchloric and nitric acid to remove any graphitic or organic residues from the surface. The NV is about $d=(10.1\pm 0.5)$ nm deep in the diamond, which we estimated by measuring the variance of the fluctuating NMR magnetic field at its location with XY8 dynamical decoupling [57].
@@
Figure S.4: (a) Pulse scheme for the CXY8-$N$ and XY8-$N$ measurements for frequency sensing. We use a laser to prepare the electron spin of an NV center in state $|0\rangle$ by optical pumping and read out its final state. We prepare a coherent superposition state by a microwave $\pi/2$ pulse, apply the control scheme (CXY8 or XY8), and then another $\pi/2$ pulse with a phase that alternates, to map coherences back onto populations. The final population of state $|0\rangle$ is estimated from the difference of the signals from the two alternating measurements, reducing errors due to charge-state and count-rate fluctuations [51].We vary $T$ ($\tau$) for CXY8-$N$ (XY8-$N$) with fixed repetition number $N$ to detect the frequency of the AC signal. (b) Measured signal, corresponding to the population of state $|0\rangle$ of the electron spin of a single NV center, vs. the detected signal frequency $f=\frac{1}{2T}$ ($f=\frac{1}{2\tau}$) for CXY8 (XY8). We observe a dip at the frequency 170017001700\,1700kHz of the artificial signal for both CXY8-5 and XY8-5 sequences, as expected from theory. The slightly lower population with XY8 than CXY8 on resonance is expected and due to the difference in the rectified signal amplitudes of the two sequences (see text). The vertical dashed lines for all subfigures show the expected full width at half maximum (FWHM) of the dip, i.e., the inverse of the sequence duration, which is FWHM $\approx 85\,$kHz in this case. (c) We observe the expected a dip in the population at 800080008000\,8000kHz of a high frequency artificial signal, where only CXY8 is applicable. FWHM $\approx 83.3\,$kHz. (d) The measurement is the same as in (b) but with a lower artificial signal amplitude, which requires a higher sequence order and total sequence duration to observe a similar dip; FWHM $\approx 16.3\,$kHz. The additional sideband dip with XY8 does not correspond to the frequency of a real signal and is likely due to pulse errors of the non-instantaneous rectangular pulses. (e) The measurement is same as in (d) but with a lower artificial signal amplitude; FWHM $\approx 16.4\,$kHz due to the slightly shorter sequence duration than in (d). All experimental results show excellent correspondence to theory.
@@
All MW pulses for the experimental sequences were generated in direct synthesis by a Keysight (M8195A) and Tektronix (AWG70001A) arbitrary waveform generators, using a Python based in house written software (Qudi) [47]. The measurement sequences are programmed to keep the phase relation between all pulses in the rotating frame and the high sampling rate of the AWGs (50 and 60 GS/s) allows to have precise direct control of the signal phases. The instantaneous rise time of the devices allows for a bandwidth as the inverse of twice the sampling rate.
@@
In a first series of experiments, we use a single NV center to detect the frequency of an artificially generated AC signal at frequency 170017001700\,1700 kHz (800080008000\,8000 kHz), which we label low (high) frequency for two different signal amplitudes. Figure S.4(a) shows the pulse scheme for the CXY8-$N$ and XY8-$N$ measurements. We use a laser to prepare the electron spin of the NV center in state $|0\rangle$ by optical pumping and read out its final state. We prepare a coherent superposition state by a microwave $\pi/2$ pulse, apply the control scheme (CXY8 or XY8), and then use another $\pi/2$ pulse with a phase that alternates between 00 and $180^{\circ}$, to map coherences back onto populations. We estimate the final population of state $|0\rangle$ from the difference between the signals, reducing errors due to charge-state and count-rate fluctuations [51]. All control fields originate from the same arbitrary waveform generator and amplifier. We generate the artificial external signal to be sensed from a second microwave source, so the phase of the signal at the beginning of the interaction is not fixed with respect to the control fields. We vary the interval of phase changes $T$ (the interpulse duration $\tau$) for CXY8 (XY8) to detect the frequency of the AC signal, using the relation $f=\frac{1}{2T}$ ($f=\frac{1}{2\tau}$).
Figure S.4(b) demonstrates frequency sensing of a relatively low frequency (1700 kHz) AC signal, which is typical for pulsed DD applications for NV centers. As expected from theory, the measured signal shows a dip in the population of state $|0\rangle$ at the frequency 170017001700\,1700kHz of the artificial signal, for both CXY8-5 and XY8-5 sequences. The slightly lower population with XY8 than CXY8 on resonance is expected from theory and is due to the corresponding difference in the rectified signal amplitudes of the two sequences, which are $g/2$ ($2g/\pi$) for CXY8 (XY8) with $g$ – the amplitude of the artificial signal in angular frequency units (see the main text). The expected full width at half maximum (FWHM) of the dip also matches theory and is approximately equal to the inverse of the sequence duration $8N\times\frac{1}{2f}$ ms on resonance. Thus, FWHM $\approx\frac{2f}{8N}=\frac{2\times 1700}{8\times 5}=85\,$kHz for both CXY8 and XY8.
Figure S.4(c) demonstrates the application of CXY8-24 for detecting a high frequency AC field, where standard pulsed DD, e.g., with the XY8 sequence, cannot be applied. We observe the expected dip in the population at the frequency 800080008000\,8000kHz of the artificial signal with a FWHM $\approx\frac{2f}{8N}=\frac{2\times 8000}{8\times 24}=83.3\,$kHz.
@@
We also perform quantum heterodyne (Qdyne) measurements with the single NV center and an 8 MHz test signal. For details on the measurement technique see the main text and references [41, 61, 48].
@@
Figure 4 in the main text shows the measured spectrum of the Qdyne measurement. In Figure S.6 we show further analysis of the performance of the Qdyne measurement. The analysis is done equivalently to Ref. [61]. The total measurement time is $T_{\mathrm{tot}}=120$ s. For the analysis we slice the data into smaller parts. The shortest time is 0.57 s (216 samples). For every measurement time the FFT of the corresponding time trace is calculated. A Lorentzian fit of the form